What is the present value of $85,001 on December 6, 2023 using 6.69%? This is the same question asked a different way. How mu$54,321 $62,255 $65,604 $79,671
Answer #1

PV = FV * {1/( 1+r)} = $ 85001 *{ 1/(1+0.0669)* } =$ 85001 * (1/1.2956) =$85001 * 0.7718 =$ 65604(approx) Answer is Option 3